4-3(2b+4) = 6-2(3b-1)
Answer by checkley75(3666)

4-3(2B+4)=6-2(3B-1)
4-6B+12=6-6B+2
-6B+6B=6+2-4-12
0=-8 NO SOLUTION FOR B.